# A Spectroscopic Search for Dormant Black Holes in Low-Metallicity Binaries

### v1

@@ 1 Introduction
2 Motivation
We design our search to be sensitive to Gaia BH3-like binaries. We use the gaiamock code (K. El-Badry et al., 2024), which forward-models Gaia epoch astrometry and model fitting, to simulate epoch astrometry and a radial velocity (RV) curve for Gaia BH3 as it would have been observed in Gaia DR3. We show the epoch astrometry and RV curve in Figure 1, with times provided relative to the DR3 reference epoch of Julian Date (JD) . The epoch astrometry was insufficiently constraining for Gaia BH3 to receive an orbital solution in DR3. However, the source had RUWE , indicative of a poor single-star astrometric fit (L. Lindegren, 2018). In addition, the star’s RV is predicted to be different from the reported mean RV in DR3 by km s-1 over of its orbit. This suggests that a search strategy based on measuring epoch RVs for low-metallicity stars with high RUWE values in DR3, and comparing those RVs to the corresponding mean RVs reported in DR3, could lead to the discovery of BH or NS binaries that did not receive orbital solutions in DR3.
@@ 2 Motivation
In general, giants lead to slightly larger RUWE values than dwarfs. However, the behavior of the curves in Figure 2 is determined by how the epoch astrometry uncertainties vary with apparent -band magnitude. The uncertainties are smallest at –, and rise at brighter or fainter magnitudes outside this range (B. Holl et al., 2023). This leads to a somewhat non-intuitive dependence of RUWE on distance in Figure 2. As sources become too bright, uncertainties rise and RUWE falls, causing giants to have smaller RUWE values than dwarfs at close distances. Companions to giants are more detectable than companions to dwarfs only at large distances ( kpc), because the uncertainties for dwarfs increase rapidly at those distances.
Our main sample includes a selection cut of RUWE (see Section 3 for details). This implies that we are sensitive to orbital periods of – d and distances out to kpc.
@@ 3 Sample Selection
We measured at least one RV for 423 of the remaining 456 candidates. We have thus observed % of the remaining 456 objects in the sample, or % of 595 when we include the objects with orbital solutions. Of the 456 candidates lacking orbital solutions, 108 have acceleration or variable acceleration solutions reported in the gaiadr3.nss_acceleration_astro catalog. 102 of these sources are in our observed sample. The acceleration solutions suggest that these are long-period ( d; K. El-Badry et al. 2024) binaries, but are insufficient to constrain the companion masses on their own.
@@ 3.3 Summary of the sample
DR3 parallax uncertainties for sources with elevated RUWE are underestimated by a factor of – (K. El-Badry, 2025). After inflating the parallax uncertainties of our targets based on the prescription of K. El-Badry (2025), we find that still have parallax to parallax uncertainty ratios , implying that this effect does not significantly affect our sample selection or our subset of BH binary candidates displaying large RV variability.
@@ 4.1 Spectroscopic Follow-up
We obtain 648 spectra for 526 targets, years after the median epoch of the time series data in Gaia DR3. We use a typical exposure time of 1800s for targets with , and shorter and longer exposure times for brighter and fainter objects.
@@ 4.1.1 FEROS
We use the Fiberfed Extended Range Optical Spectrograph (FEROS; A. Kaufer et al., 1999) on the 2.2m ESO/MPG telescope at La Silla Observatory for targets at (programs 113.26XB.001, 114.27SS.001, and 115.28KE.001). The resulting spectra have resolution over a spectral range of – nm, and a typical SNR of 4 at 5178 Å. We reduced the data using the CERES pipeline (R. Brahm et al., 2017). The pipeline performs bias subtraction, flat fielding, wavelength calibration, and optimal extraction, and measures and corrects for small shifts in the wavelength solution over the course of a night using simultaneous observations of a ThAr lamp with a second fiber. For each spectrum, we estimate the continuum using a median filter, and then perform continuum normalization.
@@ 4.1.2 APF
We use the Levy spectrometer on the 2.4m Automated Planet Finder (APF; M. V. Radovan et al., 2010; S. S. Vogt et al., 2014) telescope at Lick Observatory for targets at . We used a 2” slit, yielding spectra with resolution over a wavelength range of – nm and a typical SNR of 8 at 5178 Å. We reduced the data using the California Planet Search pipeline (A. W. Howard et al., 2010; B. J. Fulton et al., 2015). As with the FEROS spectra, we estimate the continuum of each spectrum using a median filter, and then perform continuum normalization.
@@ 4.2 Radial Velocities
We measure RVs via cross-correlating with a synthetic template spectrum. Specifically, we measure RVs based on two consecutive orders that display prominent absorption features. We use orders 9 and 10 for FEROS and orders 34 and 35 for APF, corresponding to wavelength regions of 5048–5362 Å and 5137–5279 Å, respectively. These wavelength regions include the Mg I b triplet, which in most of our targets is among the strongest features in the optical. We experimented with a variety of spectral regions and found these to be the most robust; the Balmer lines are too broad, and the Mg I lines are among the strongest lines in metal-poor stars. Many other orders do not contain detectable lines in most of our targets.
@@
For each target, we choose the template spectrum as the BOSZ Kurucz (R. L. Kurucz, 1979, 1993; S. Mészáros et al., 2024) spectrum with stellar parameters closest to those derived by R. Andrae et al. (2023). The template spectra cover a grid of metallicity spaced by , effective temperature spaced by K, and surface gravity spaced by . The template spectra are at a resolution of and assume no rotational broadening.
@@
We manually inspected the template fit in all cases where the epoch RV differed from the mean DR3 RV by more than 10 km s-1 (see Table 1). During this vetting process, we visually identified a few objects where a joint fit to both orders produced a less accurate RV than a fit to only one of the orders. In these cases, we adopt the more accurate RV derived from the best fit to one of the orders alone.
@@ 5 Results
If we account for the uncertainties on our RV measurements and the reported Gaia DR3 RVs, only 48 of the 66 interesting candidates listed in Table 1 have RV discrepancies at the level. However, these uncertainties are dominated by the uncertainty in the Gaia DR3 mean RV, which is often a result of orbital motion rather than the statistical uncertainty on an epoch RV measurement (see Section 6 for further details on the process by which DR3 RVs are computed). As such, we still include the 18 candidates whose largest RV shift is statistically consistent with zero to within , as we do not think that this substantially undermines the evidence indicating the presence of a potential compact object companion.
@@ 5.1 Objects with Multiple RVs
We attempted to obtain at least one additional follow-up RV for each target with RV discrepancy km s-1. We plot RVs over time for two typical targets in Table 1 with RV discrepancy km s-1 in Figure 7. The target in the left panel displays RV variability on short timescales. This suggests that the source is an unresolved hierarchical triple; in this scenario, the observed RV variability arises from the putative inner binary, while the large RUWE value is due to the motion of the outer tertiary. The target in the right panel displays slower RV variability that suggests a long orbital period, consistent with our expectations for typical binaries in our sample.
While the total number of RV epochs per target ranges from –, and we observe some RV variability, we have not yet fully solved the orbit of any of our targets. For targets with orbital periods days, a few RVs taken over a short time frame are insufficient to fit a full orbital solution and tightly constrain the companion mass. Since extended RV monitoring is necessary to determine the orbital period and confirm or rule out a BH or NS companion, we will continue following up interesting candidates over the next several months. Upon release of DR4, we will derive constraints on the companion masses by jointly fitting our long-baseline measurements with epoch Gaia astrometry.
@@ 5.2 A Promising Candidate
Using the approach described by H. Reggiani et al. (2022), we jointly fit the SED and high-resolution spectrum to derive atmospheric parameters for this target. We also measure the equivalent widths of atomic absorption lines to infer elemental abundances. We provide our best-fit parameters and abundances in Table 2. For comparison, R. Andrae et al. (2023) estimate , K, and . The atmospheric parameters we derive are typical for a low-metallicity halo star (given our selection criteria), and are in reasonable agreement with those provided by R. Andrae et al. (2023). The target’s abundance pattern is unremarkable for a low-metallicity halo star. We verify that non-local thermodynamic equilibrium corrections do not strongly affect any of our derived abundances.
@@ 6.2 How many low-metallicity stars in our sample have BH companions?
To estimate the answer, we simulate a population of BH + luminous star binaries in the Galactic halo. We draw stars from a Kroupa initial mass function (IMF) (P. Kroupa, 2001) between and . We sample Galactocentric radii from a power-law distribution, assuming a halo density profile (e.g., X.-X. Xue et al. 2015; C. T. Slater et al. 2016; A. J. Deason et al. 2018; studies that adopt a double power-law profile find a shallower inner slope and steeper outer slope). We then calculate heliocentric distances and retain the stars that lie kpc away from the Sun, which we place at a distance of 8.122 kpc away from the Galactic Center (GRAVITY Collaboration et al., 2018) and at a height of 20.8 pc above the Galactic mid-plane (M. Bennett & J. Bovy, 2019). We estimate absolute magnitudes using synthetic photometry from a 12 Gyr MIST isochrone with initial metallicity , and use them to calculate apparent -band magnitudes. We draw orbital periods from a log-uniform distribution between and days (reasonable for Gaia BH3-like systems formed in clusters, see e.g., D. Marín Pina et al. 2024) and eccentricities from a thermal distribution. We assign sky locations by randomly sampling right ascensions and declinations from our catalog of candidate metal-poor binaries. Then, assuming that each luminous star has a BH companion, we randomly sample binary orientations and calculate DR3 RUWE values using the gaiamock code (K. El-Badry et al., 2024). Finally, we apply selection cuts of RUWE and to identify the simulated sources that would make it into our sample. We display the properties of our simulated sample of targets in Figure 10, confirming that brighter and closer binaries tend to have larger predicted RUWE values in DR3.
Based on the location of the Hertzsprung Gap in the color-magnitude diagram in the left panel of Figure 10, we now define giants as stars with absolute -band magnitude and dwarfs as stars with . We find that there are times more dwarf-BH systems than giant-BH systems in the simulated low-metallicity sample; while the IMF predicts that there are more than 8 dwarfs for every giant, most of those dwarfs are too faint to make it into the sample. Thus, if Gaia BH3 is the only giant-BH system in our observed sample, we expect there to be undiscovered dwarf-BH systems in our observed sample as well. As discussed in Section 6.3 below, we expect our search strategy to recover about half of these compact object binaries.
@@ 6.3 Search sensitivity
Using gaiamock, we compute the fraction of mock observed BH binaries that have both predicted DR3 RUWE and a discrepancy between their epoch RV and mean DR3 RV greater than our adopted threshold of km s-1. To derive the uncertainty on this fraction, we repeat our simulation times. We plot our results in Figure 11. We find that, assuming typical orbital periods and BH masses, our search strategy recovers about half of BHs in low-metallicity binaries that lie closer than kpc and have luminous companions brighter than . Some real BH binaries may not be recovered if they have orientations close to face-on or are observed at a time during which their epoch RV is similar to their average RV. The search sensitivity peaks at an orbital period of d, with the sensitivity being significantly lower at much shorter or longer orbital periods.
@@ 7 Conclusion
Summary of the sample: We select a sample of bright, metal-poor sources with and excess astrometric noise in DR3, adopting the robust, data-driven metallicities derived from XP spectra by R. Andrae et al. (2023). We use two different selection cuts (Section 3) for targets with and RUWE and targets with , RUWE , and large estimated , respectively. We apply additional quality cuts to exclude marginally resolved binaries and stars with disk-like orbits. We obtained at least one FEROS or APF spectrum for 526 targets out of a sample of 717 candidates, with the remainder scheduled to be observed in the future. The sample of observed targets has , RUWE , –, and is bluer than the main sequence on a color-magnitude diagram (Figure 3).

### v2

@@ 1 Introduction
2 Search Strategy
We design our search to be sensitive to Gaia BH3-like binaries. We use the gaiamock code (K. El-Badry et al., 2024), which forward-models Gaia epoch astrometry and model fitting, to simulate epoch astrometry and a radial velocity (RV) curve for Gaia BH3 as it would have been observed in Gaia DR3. We show the epoch astrometry and RV curve in Figure 1, with times provided relative to the DR3 reference epoch of Julian Date (JD) . The epoch astrometry was insufficiently constraining for Gaia BH3 to receive an orbital solution in DR3. However, the source had RUWE , indicative of a poor single-star astrometric fit (L. Lindegren, 2018). In addition, the star’s RV is predicted to be different from the reported mean RV in DR3 by km s-1 over of its orbit. This suggests that a search strategy based on measuring epoch RVs for low-metallicity stars with high RUWE values in DR3, and comparing those RVs to the corresponding mean RVs reported in DR3, could lead to the discovery of BH or NS binaries that did not receive orbital solutions in DR3. Indeed, RV follow-up complements and validates RUWE-based filtering, with even a few RVs being able to break the degeneracy between low companion mass, long-period systems and high companion mass, short-period systems that both have elevated RUWE. The interpretation of any observed RV discrepancy between a binary’s epoch RVs and its mean DR3 RV depends on its true orbital period (see Section 6.3).
@@ 2 Search Strategy
In general, giants lead to slightly larger RUWE values than dwarfs, though there is a somewhat non-intuitive dependence of RUWE on distance in Figure 2. This behavior is determined by how the epoch astrometry uncertainties vary with apparent -band magnitude. The uncertainties are smallest at –, and rise at brighter or fainter magnitudes outside this range (B. Holl et al., 2023). As sources become too bright, uncertainties rise and RUWE falls, causing giants to have smaller RUWE values than dwarfs at close distances. Companions to giants are more detectable than companions to dwarfs only at large distances ( kpc), because the uncertainties for dwarfs increase rapidly at those distances.
Our main sample includes a selection cut of RUWE (see Section 3 for details). This implies that we are sensitive to orbital periods of – d and distances out to kpc. That being said, binaries that have elevated RUWE can show no detectable RV discrepancy if they have high eccentricities (depending on DR3 phase coverage) or face-on orbits. We quantify this in Section 6.
@@ 3 Sample Selection
We measured at least one RV for 425 of the remaining 456 candidates. We have thus observed % of the remaining 456 objects in the sample, or % of 595 when we include the objects with orbital solutions. Of the 456 candidates lacking orbital solutions, 108 have acceleration or variable acceleration solutions reported in the gaiadr3.nss_acceleration_astro catalog. 102 of these sources are in our observed sample. The acceleration solutions suggest that these are long-period ( d; K. El-Badry et al. 2024) binaries, but are insufficient to constrain the companion masses on their own.
@@ 3.3 Summary of the sample
Our selection cuts depend on the availability of robustly measured parallaxes, but DR3 parallax uncertainties for sources with elevated RUWE are underestimated by a factor of – (K. El-Badry, 2025). After inflating the parallax uncertainties of our targets based on the prescription of K. El-Badry (2025), we find that still have parallax to parallax uncertainty ratios , implying that most of our candidates still have reasonably well-constraned distances.
@@ 4.1 Spectroscopic Follow-up
We obtained 657 spectra for 528 targets, years after the median epoch of the time series data in Gaia DR3. Of these spectra, 499 were obtained using FEROS, and 158 were obtained using APF. We used a typical exposure time of 1800s for targets with , and shorter and longer exposure times for brighter and fainter objects. We make all our spectra publicly available at https://doi.org/10.5281/zenodo.16888415.
@@ 4.1.1 FEROS
We used the Fiberfed Extended Range Optical Spectrograph (FEROS; A. Kaufer et al., 1999) on the 2.2m ESO/MPG telescope at La Silla Observatory for targets at (programs 113.26XB.001, 114.27SS.001, and 115.28KE.001). The resulting spectra have resolution over a spectral range of – nm, and a typical SNR of 4 at 5178 Å. We reduced the data using the CERES pipeline (R. Brahm et al., 2017). The pipeline performs bias subtraction, flat fielding, wavelength calibration, and optimal extraction, and measures and corrects for small shifts in the wavelength solution over the course of a night using simultaneous observations of a ThAr lamp with a second fiber. For each spectrum, we estimated the continuum using a median filter, and then performed continuum normalization.
@@ 4.1.2 APF
We used the Levy spectrometer on the 2.4m Automated Planet Finder (APF; M. V. Radovan et al., 2010; S. S. Vogt et al., 2014) telescope at Lick Observatory for targets at . We used a 2” slit, yielding spectra with resolution over a wavelength range of – nm and a typical SNR of 8 at 5178 Å. We reduced the data using the California Planet Search pipeline (A. W. Howard et al., 2010; B. J. Fulton et al., 2015). We measured and corrected for small shifts in the wavelength solution over the course of a night using telluric absorption lines (see P. Nagarajan et al. 2023 for a description of the method). As with the FEROS spectra, we estimated the continuum of each spectrum using a median filter, and then performed continuum normalization.
@@ 4.2 Radial Velocities
We measured RVs via comparison with a synthetic template spectrum. Specifically, we measured RVs based on two consecutive orders that display prominent absorption features. We used orders 9 and 10 for FEROS and orders 34 and 35 for APF, corresponding to wavelength regions of 5048–5362 Å and 5137–5279 Å, respectively. These wavelength regions include the Mg I b triplet, which in most of our targets is among the strongest features in the optical. We experimented with a variety of spectral regions and found these to be the most robust; the Balmer lines are too broad, and the Mg I lines are among the strongest lines in metal-poor stars. Many other orders do not contain detectable lines in most of our targets.
@@
For each target, we chose the template spectrum as the BOSZ Kurucz (R. L. Kurucz, 1979, 1993; S. Mészáros et al., 2024) spectrum with stellar parameters closest to those derived by R. Andrae et al. (2023). The template spectra cover a grid of metallicity spaced by , effective temperature spaced by K, and surface gravity spaced by . The template spectra are at a resolution of and assume no rotational broadening.
@@
We manually inspected the template fit in all cases where the epoch RV differed from the mean DR3 RV by more than 10 km s-1 (see Table 1). We do not account for the error in the mean DR3 RV when adopting this threshold, since that error often arises from orbital motion rather than statistical uncertainty. During this vetting process, we visually identified a few objects where a joint fit to both orders produced a less accurate RV than a fit to only one of the orders. In these cases, we adopt the more accurate RV derived from the best fit to one of the orders alone.
@@ 5 Results
If we account for the uncertainties on our RV measurements and the reported Gaia DR3 RVs, only 49 of the 67 interesting candidates listed in Table 1 have RV discrepancies at the level. However, these uncertainties are dominated by the uncertainty in the Gaia DR3 mean RV, which is often a result of orbital motion rather than the statistical uncertainty on an epoch RV measurement (see Section 6 for further details on the process by which DR3 RVs are computed). As such, we still include the 18 candidates whose largest RV shift is statistically consistent with zero to within , as we do not think that this substantially undermines the evidence indicating the presence of a potential compact object companion.
@@ 5.1 Objects with Multiple RVs
We are attempting to obtain at least one additional follow-up RV for each target with RV discrepancy km s-1. We plot RVs over time for two typical targets in Table 1 with RV discrepancy km s-1 in Figure 7. The target in the left panel displays RV variability on short timescales. This suggests that the source is an unresolved hierarchical triple; in this scenario, the observed RV variability arises from the inner binary, while the large RUWE value is due to the motion of the outer orbit. The target in the right panel displays slower RV variability that suggests a long orbital period, consistent with our expectations for typical binaries in our sample.
For one target in our sample, Gaia DR3 2123589958955565824, we measured 11 RVs, which is enough to constrain the full orbital solution. We use ensemble MCMC sampling (emcee; D. Foreman-Mackey et al., 2013) with 64 walkers and 10000 iterations to fit the RVs with a two-body Keplerian orbit and show the best-fit RV curve in Figure 8. We find a short orbital period of d and a moderate eccentricity of . The best-fit RV semi-amplitude of km s-1 leads to a mass function of and a minimum companion mass of , assuming a luminous star mass of . This makes the system unlikely to contain a BH. Using gaiamock, we find that the predicted DR3 RUWE for the derived orbit and typical inclinations is , much smaller than the observed DR3 RUWE of . We conclude that the system is likely a hierarchical triple, with the RVs tracing the motion of the inner binary.
While the total number of RV epochs per target ranges from –, and we observe some RV variability in all targets, we have not yet fully solved the orbit of any of our other targets. For targets with orbital periods days, a few RVs taken over a short time frame are insufficient to fit a full orbital solution and tightly constrain the companion mass. Since extended RV monitoring is necessary to determine the orbital period and confirm or rule out a BH or NS companion, we will continue following up interesting candidates over the next several months. Upon release of DR4, we will derive constraints on the companion masses by jointly fitting our long-baseline measurements with epoch Gaia astrometry.
@@ 5.2 A Promising Candidate
Using the approach described by H. Reggiani et al. (2022), we jointly fit the SED and high-resolution spectrum to derive atmospheric parameters for this target. We also measure the equivalent widths of atomic absorption lines to infer elemental abundances. We provide our best-fit parameters and abundances in Table 2. The reported uncertainties are statistical, and do not include systematic errors. For comparison, R. Andrae et al. (2023) estimate , K, and . The atmospheric parameters we derive are typical for a low-metallicity halo star (given our selection criteria), and are in reasonable agreement with those provided by R. Andrae et al. (2023). The target’s abundance pattern is unremarkable for a low-metallicity halo star. We verify that non-local thermodynamic equilibrium corrections do not strongly affect any of our derived abundances.
@@ 6.2 How many low-metallicity stars in our sample have BH companions?
To estimate the answer, we simulate a population of BH + luminous star binaries in the Galactic halo. We draw stars from a Kroupa initial mass function (IMF) (P. Kroupa, 2001) between and . We sample Galactocentric radii from a power-law distribution, assuming a halo density profile (e.g., X.-X. Xue et al. 2015; C. T. Slater et al. 2016; A. J. Deason et al. 2018; studies that adopt a double power-law profile find a shallower inner slope and steeper outer slope). We then calculate heliocentric distances and retain the stars that lie kpc away from the Sun, which we place at a distance of 8.122 kpc away from the Galactic Center (GRAVITY Collaboration et al., 2018) and at a height of 20.8 pc above the Galactic mid-plane (M. Bennett & J. Bovy, 2019). We estimate absolute magnitudes using synthetic photometry from a 12 Gyr MIST isochrone with initial metallicity , and use them to calculate apparent -band magnitudes. We assume zero extinction. We draw orbital periods from a log-uniform distribution between and days (reasonable for Gaia BH3-like systems formed in clusters, see e.g., D. Mar´ın Pina et al. 2024) and eccentricities from a thermal distribution. We assign sky locations by randomly sampling right ascensions and declinations from our catalog of candidate metal-poor binaries. Then, assuming that each luminous star has a BH companion, we randomly sample binary orientations and calculate DR3 RUWE values using the gaiamock code (K. El-Badry et al., 2024). Finally, we apply selection cuts of RUWE and to identify the simulated sources that would make it into our sample. We display the properties of our simulated sample of targets in Figure 11, confirming that brighter and closer binaries tend to have larger predicted RUWE values in DR3.
Based on the location of the Hertzsprung Gap in the color-magnitude diagram in the left panel of Figure 11, we now define giants as stars with absolute -band magnitude and dwarfs as stars with . We find that there are times more dwarf-BH systems than giant-BH systems in the simulated low-metallicity sample; while the IMF predicts that there are more than 8 dwarfs for every giant, most of those dwarfs are too faint to make it into the sample. Thus, if Gaia BH3 is the only giant-BH system in our observed sample, we expect there to be undiscovered dwarf-BH systems in our observed sample as well. As discussed in Section 6.3 below, we expect our search strategy to recover about half of these compact object binaries. If there are 5 genuine BH binaries in our sample, then the estimated false positive rate of our candidate sample is , or about .
@@ 6.3 Search sensitivity
Using gaiamock, we compute the fraction of mock observed BH binaries that have both predicted DR3 RUWE and a discrepancy between their epoch RV and mean DR3 RV greater than our adopted threshold of km s-1. To derive the uncertainty on this fraction, we repeat our simulation times. We plot our results in Figure 12. We find that, assuming typical orbital periods ( d) and BH masses (, our search strategy recovers about half of BHs in low-metallicity binaries that lie closer than kpc and have luminous companions brighter than . Some real BH binaries may not be recovered if they have orientations close to face-on or are observed at a time during which their epoch RV is similar to their average RV. The search sensitivity peaks at orbital periods of a few hundred days, with the sensitivity being significantly lower at much shorter or longer orbital periods.
@@ 7 Conclusion
Summary of the sample: We select a sample of bright, metal-poor sources with and excess astrometric noise in DR3, adopting the robust, data-driven metallicities derived from XP spectra by R. Andrae et al. (2023). We use two different selection cuts (Section 3) for targets with and RUWE and targets with , RUWE , and large estimated , respectively. We apply additional quality cuts to exclude marginally resolved binaries and stars with disk-like orbits. We obtained at least one FEROS or APF spectrum for 528 targets out of a sample of 717 candidates, with the remainder scheduled to be observed in the future. The sample of observed targets has , RUWE , –, and is bluer than the main sequence on a color-magnitude diagram (Figure 3).